OpenStudy (kkutie7):

how do I solve for x? \[sin(9x)=0\]

OpenStudy (kkutie7):

no

OpenStudy (zarkon):

can you solve \(\sin(y)=0\)

OpenStudy (kkutie7):

can't you just do this \[y=cos^{-1}(0)\]

OpenStudy (kkutie7):

or I can use a unit circle cant I?

OpenStudy (zarkon):

what values of y is the sine of y zero. I'll give a couple sin(0)=0 \[ \sin(\pi)=0\]

OpenStudy (kkutie7):

ok I understand that... I'm getting throw off by the 9

OpenStudy (zarkon):

so ... can you give me all the solutions to \[\sin(y)=0\]

OpenStudy (zarkon):

then we will take care of the 9

OpenStudy (kkutie7):

ok I can do that \[y=\frac{\pi}{2},\pi,0, and \frac{3\pi}{2} \]

OpenStudy (zarkon):

pi/2 and 3pi/2 are not solutions

OpenStudy (kkutie7):

shoot you are right that is for cos

OpenStudy (zarkon):

but 2pi is 3pi, 4pi -pi -2pi

OpenStudy (kkutie7):

to be fair I'm gonna need that for this problem in a bit =)

OpenStudy (kkutie7):

\[-\frac{\pi}{27}< x < \frac{\pi}{27}\]

OpenStudy (zarkon):

so solve \(\sin(9x)=0\) when \(-\dfrac{\pi}{27}< x < \dfrac{\pi}{27}\)

OpenStudy (zarkon):

?

OpenStudy (kkutie7):

yes

OpenStudy (zarkon):

then there is only one solution

OpenStudy (zarkon):

when sin(0)=0 x=0

OpenStudy (zarkon):

the next largest solution would be \(sin(\pi)=0\) and so \(9x=\pi\) ans \(x=\dfrac{\pi}{9}\) which is outside the interval (same for -\(\pi\))

OpenStudy (kkutie7):

so x=0

OpenStudy (zarkon):

yes

OpenStudy (kkutie7):

so for cos(9x)=1/2 would I use pi/3?

OpenStudy (kkutie7):

no that too big right?

OpenStudy (zarkon):

on what interval? same?

OpenStudy (kkutie7):

yes

OpenStudy (zarkon):

well \(\cos(\pi/3)=1/2\) set \[9x=\pi/3\] so \[x=\pi/27\] is a solution but it is not in your interval

OpenStudy (zarkon):

same goes for \(-\pi/3\) so there are no solutions on \(-\dfrac{\pi}{27}< x < \dfrac{\pi}{27}\)

OpenStudy (kkutie7):

Thank you so much

OpenStudy (zarkon):

if it were on the interval \(-\dfrac{\pi}{27}\le x \le\dfrac{\pi}{27}\) then \(-\pi/27\) and \(\pi/27\) would be solutions

OpenStudy (nikato):

actually, isnt there two solutions for sin(9x)=0 ?

OpenStudy (zarkon):

what would the other solution be?

OpenStudy (nikato):

when x=20

OpenStudy (nikato):

because sin(0)=0 and also sin(180)=0

OpenStudy (zarkon):

this is in radians

OpenStudy (zarkon):

180\(^\circ\) is the same as \(\pi\)rad and \(\pi\) is not a solution

OpenStudy (zarkon):

since it is on within our interval \(-\dfrac{\pi}{27}< x < \dfrac{\pi}{27}\)

OpenStudy (zarkon):

in particular \(\pm\dfrac{\pi}{9}\) is ont in the interval

OpenStudy (nikato):

oh ok. nvm. i did not know there was a restriction